OpenStudy (anonymous):

If the volume of a sphere increases by 72.8%, then its surface area increases by how much and how ........

OpenStudy (mayankdevnani):

The area will increase by ((1+.728)^(2/3)-1)*100% can you solve it? @mondalabhilash

OpenStudy (anonymous):

\[Vol \propto r^3\]\[SA\propto r^2\]

OpenStudy (anonymous):

\[V'=1.728V\] \[r'^3=1.728r^3\] \[(\frac{r'}{r})^3=1.728\] Now find \[(\frac{r'}{r})^2\], and rearrange to get \[r'^2=???r^2\]